Week 40: September 28 - October 5 Last updated:   29-Sep-20 17:15
Maintenance this week Date Time & Length Production resumes
520 MeV Tuesday, Sept. 29 maintenance from 08:30 - 14:30 (6 hr), followed by 2 hours start-up. 16:30
ISAC Tuesday, Sept. 29 maintenance starting at 08:30
E-LINAC Maintenance on-going, E-Hall open all week: isolation transformer + beam dump work
ARIEL
                   
GROUP ACTIVITY Fault Report Work Permit Completed
ATG Crane tgt flask from ATG -> STF hotcell 520 MeV C2020-09-29-11 y
Water package work on SAB3 520 MeV - n
Beamlines
Computing and Networking
Controls Test SSH server restart in DEVBN2 520 MeV - n
New cooling tower controls work 520 MeV - on-going
Test water systems TCs 520 MeV - n
STF Break-out Panel Reconfiguration  520 MeV - y
ISAC Target Hall TM4 move from the TCS to ITE followed by TM4 service connections. ISAC I2020-09-29-4, I2020-09-29-2 y
ISIS RIB mock assembly move to RH Hot Cell in MHE 520 MeV C2020-09-29-4 y
Check I1 H2 supply  520 MeV - n
Operations Sump checks All - y
SCBCD & SCRM Tests 520 MeV - y
Remote Handling
RF Take measurements inside the PAs for test stand project (RF off ~ 30  min) 520 MeV C2020-09-29-1
Safety Replace RMS1 PC on SAB2 (beam spill monitors will go down ~10 min) 520 MeV C2020-09-29-3 y
Site Services SAB3 pump check 520 MeV - n
Mechanical, Civil/Structural, Electrical, FAS Connect control cables to old cooling tower vfds (may need RF off, tbd) 520 MeV C2020-09-29-5 y
Craning - RIB mock assembly move to RH Hot Cell in MHE 520 MeV C2020-09-29-4 y
Hazmat sampling in CMMS B2 storage area (needs 1A off) 520 MeV A2020-09-29-3 y
Inspect and troubleshoot subsurface sump pumps 520 MeV 13444 C2020-09-29-8 y
Swap out dissolved oxygen senor on Cu-ALCW side stream 520 MeV 13448 C2020-09-29-7 To be continued
E-LINAC HAW flush in SAB3 Tunnel E-LINAC/520 MeV A2020-09-29-3 y
Add thermal overload unit to Cu-ALCW deionizer loop pump 520 MeV C2020-09-29-9 y